# Can KAN Work? Exploring the Potential of Kolmogorov-Arnold Networks in Computer Vision

### v1

@@ Abstract
Kolmogorov-Arnold Networks (KANs), as a theoretically efficient neural network architecture, have garnered attention for their potential in capturing complex patterns. However, their application in computer vision remains relatively unexplored. This study first analyzes the potential of KAN in computer vision tasks, evaluating the performance of KAN and its convolutional variants in image classification and semantic segmentation. The focus is placed on examining their characteristics across varying data scales and noise levels. Results indicate that while KAN exhibits stronger fitting capabilities, it is highly sensitive to noise, limiting its robustness. To address this challenge, we propose a smoothness regularization method and introduce a Segment Deactivation technique. Both approaches enhance KAN’s stability and generalization, demonstrating its potential in handling complex visual data tasks.
@@ 1 Introduction
In recent years, neural network[11, 9, 19] architectures have achieved remarkable breakthroughs in computer vision, especially in tasks such as image classification, object detection, and semantic segmentation[17, 22, 23]. Traditional models, particularly Convolutional Neural Networks (CNNs)[18, 17, 28, 10, 13] and the more recent Transformer architectures[33, 4, 6, 4, 20, 34, 35, 5], have been widely adopted in these areas due to their ability to capture spatial hierarchies and long-range dependencies. CNNs leverage translational invariance through convolutional filters, while Transformers effectively handle complex patterns in images using self-attention mechanisms. These models have thus become mainstream approaches in computer vision tasks.
Alongside these established models, Kolmogorov-Arnold Networks (KANs)[21] have emerged as an intriguing alternative due to their theoretical efficiency and compact design. KANs are inspired by the Kolmogorov-Arnold representation theorem[14, 15], which suggests that any continuous multivariate function can be represented by a sum of univariate functions. Based on this principle, KANs utilize learnable activation functions, such as B-spline functions[29, 27], allowing them to adapt flexibly to complex patterns with fewer parameters compared to traditional networks. This characteristic makes KANs theoretically appealing for applications where model interpretability and parameter efficiency are prioritized, thus attracting significant research interest[32, 8, 2, 1, 3].
This study aims to assess the potential and limitations of KANs in vision tasks by systematically evaluating KANs and their variations across several core computer vision tasks, including image classification, object detection, and semantic segmentation. In our experiments, we examined the response of KAN and Convolutional KAN[1] models under varying data scales and noise levels. The results demonstrate that while KAN and Convolutional KAN exhibit stronger fitting capabilities with larger datasets, their high sensitivity to noise limits their robustness in vision applications.
@@
1. Evaluation of KANs in Vision Tasks. We assess Kolmogorov-Arnold Networks (KANs) across key computer vision tasks, including image classification, object detection, and segmentation, identifying both their strengths and limitations in high-dimensional contexts.
@@ 4 Experiments
In this section, we present a series of experiments designed to evaluate the performance of Kolmogorov-Arnold Networks (KAN) and Convolutional KAN (CKAN) across several core computer vision tasks. These preliminary experiments help us assess the strengths and limitations of KAN in visual pattern recognition, providing a foundation for more in-depth analysis. Specifically, we apply KAN and CKAN to image classification and semantic segmentation tasks, using three widely-used datasets: CIFAR-10 [16], CIFAR-100 [16], and PASCAL VOC2012 [7]. Our goal is to provide a representative evaluation of KAN’s capabilities and limitations in computer vision through comprehensive experimentation. Subsequently, we will conduct an in-depth analysis and discussion based on the experimental results and consider potential ways to improve KAN’s performance in computer vision tasks.
@@ 4.1 Experiment 0: Preliminary Evaluation of KAN and CKAN on Core Vision Tasks
Result. As shown in Table 1, the experimental results indicate that the performance of models incorporating KAN and CKAN in classification and segmentation tasks is generally lower than that of the baseline models. This result is unexpected, as the nonlinear characteristics of KAN and CKAN could theoretically enhance the model’s fitting capability. This observation motivates further investigation into why the nonlinear fitting abilities of KAN and CKAN have not been fully leveraged in computer vision tasks.
@@
MobileNet First Layer CKAN 59.76
MobileNet Last Layer CKAN 60.29
@@
UNet CKAN (Final CNN replaced) 62.13
@@ 4.2 Experiment 1: Effect of Dataset Size on KAN Performance
Results. As shown in Table 3, CNN + KAN consistently outperformed the baseline CNN + MLP across all dataset sizes, demonstrating the stable fitting capability of KAN. CKAN + CNN + MLP, however, only surpassed the baseline at the 100% dataset size, indicating that CKAN’s fitting ability becomes more evident with larger data volumes. These findings reveal that both KAN and CKAN possess stronger nonlinear fitting capabilities compared to the baseline. However, as model complexity and dataset size increase, their fitting potential can be more fully realized. This experiment validates the enhanced capacity of KAN and CKAN to capture complex patterns, especially with sufficient data.
@@
4.3 Experiment 2: Robustness of KAN and CKAN to Label Noise
In Experiment 1, we observed that both KAN and CKAN exhibit strong fitting capabilities. However, theoretically, CKAN should have a superior fitting ability compared to KAN. Nevertheless, in our experimental results, CKAN’s performance was less than satisfactory. We hypothesize that this may be due to the higher complexity of input in vision tasks, which are more likely to include noise and outliers. This increased fitting capacity may, therefore, lead to lower accuracy on the test set. To investigate this, we designed Experiment 2, where we intentionally introduce label noise to evaluate the performance of CKAN and KAN in noisy environments, aiming to assess their robustness against noise in vision tasks.
@@ 4.3 Experiment 2: Robustness of KAN and CKAN to Label Noise
Results. As shown in Table 4, the accuracy of all models decreases as noise levels increase, with CKAN+CNN+MLP and CNN+KAN experiencing a more pronounced decline compared to the baseline CNN model. Although the initial performance of KAN and CKAN models is slightly better than the baseline under low noise conditions, they exhibit greater sensitivity to noise at higher noise levels. This suggests that, despite the strong fitting capabilities of KAN and CKAN, these abilities may lead to overfitting when noise is introduced, thus reducing their robustness in noisy environments.
@@ 4.4 Experiment 3: Impact of Regularization on the Generalization of KAN and CKAN
Results. As shown in Tables 5 and 6, we observed that with moderate L1 regularization, models incorporating KAN and CKAN surpassed the baseline CNN model in accuracy. This suggests that KAN and CKAN’s fitting capabilities were better utilized without being overly affected by noise points and outliers, highlighting the importance of regularization for enhancing their generalization. Additionally, we observed that while CKAN demonstrated strong fitting ability, its performance was generally inferior to that of KAN, with CKAN also requiring more memory and being computationally slower. Based on these findings, we focus on KAN in the next experiments to further explore its potential.
@@ 4.5 Experiment 4: Enhanced Regularization and Segment Deactivation for Improved KAN Generalization
MobileNet + KAN + Smoothness Regularization Method + Segment Deactivation Technique 62.98
@@ 5 Conclusions
This study explores the potential of Kolmogorov-Arnold Networks (KAN) in computer vision tasks. Despite KAN’s strong fitting capability, our experimental results indicate that the high complexity of visual data and the prevalence of noise present significant challenges to its generalization ability. To address these issues, we introduced a smoothness regularization method and the Segment Deactivation technique, effectively enhancing KAN’s robustness by reducing overfitting to noise and promoting model smoothness.

### v2

# KAN not Work: Investigating the Applicability of Kolmogorov-Arnold Networks in Computer Vision

@@ Abstract
Kolmogorov-Arnold Networks (KAN) have gained attention for their potential in capturing complex patterns, but the applicability of KAN and its variant, Convolutional KAN (CKAN), in computer vision remains a subject of significant debate. This study analyzes the performance of KAN and CKAN in computer vision tasks, with a focus on their robustness across different data scales and noise levels. The results show that CKAN, due to its extremely low generalization ability, often performs worse than traditional CNNs. Although the original KAN demonstrates strong fitting capabilities, its generalization ability is also weak, making it unsuitable for replacing the final layer of a model with an NLP layer. To address this issue, we propose a smoothness regularization method and a ”Segment Deactivation” technique, both of which significantly improve the generalization ability of KAN, leading to enhanced performance when replacing the final NLP layer.
@@ 1 Introduction
In recent years, neural network[11, 9, 19] architectures have achieved remarkable breakthroughs in computer vision, especially in tasks such as image classification, and semantic segmentation[17, 22, 23]. Traditional models, particularly Convolutional Neural Networks (CNNs)[18, 17, 28, 10, 13] and the more recent Transformer architectures[33, 4, 6, 4, 20, 34, 35, 5], have been widely adopted in these areas due to their ability to capture spatial hierarchies and long-range dependencies. CNNs leverage translational invariance through convolutional filters, while Transformers effectively handle complex patterns in images using self-attention mechanisms. These models have thus become mainstream approaches in computer vision tasks.
Alongside these established models, Kolmogorov-Arnold Networks (KAN)[21] have emerged as an intriguing alternative due to their theoretical efficiency and compact design. KAN are inspired by the Kolmogorov-Arnold representation theorem[14, 15], which suggests that any continuous multivariate function can be represented by a sum of univariate functions. Based on this principle, KAN utilize learnable activation functions, such as B-spline functions[29, 27], allowing them to adapt flexibly to complex patterns with fewer parameters compared to traditional networks. This characteristic makes KAN theoretically appealing for applications where model interpretability and parameter efficiency are prioritized, thus attracting significant research interest[32, 8, 2, 1, 3].
This study aims to systematically evaluate the performance of KAN and its variants across several core computer vision tasks to explore their potential and limitations in vision applications, including image classification and semantic segmentation. In our experiments, we assessed the performance of KAN and Convolutional KAN[1] models under varying data scales and noise levels. The results show that although KAN and Convolutional KAN exhibit strong fitting capabilities theoretically, their high sensitivity to noise severely impacts their robustness in vision tasks, leading to poor performance in practical vision applications.
@@
1. Evaluation of KANs in Vision Tasks. We evaluated the performance of Kolmogorov-Arnold Networks (KAN) and its variants in key computer vision tasks, including image classification and segmentation, and identified significant generalization issues in their application to computer vision tasks.
@@ 4 Experiments
In this section, we present a series of experiments designed to evaluate the performance of Kolmogorov-Arnold Networks (KAN) and Convolutional KAN (CKAN) across several core computer vision tasks. These preliminary experiments help us assess the strengths and limitations of KAN in visual pattern recognition, providing a foundation for more in-depth analysis. Specifically, we apply KAN and CKAN to image classification and semantic segmentation tasks, using three widely-used datasets: CIFAR-10 [16], CIFAR-100 [16], and PASCAL VOC2012 [7]. Our goal is to provide an evaluation of KAN’s capabilities in computer vision tasks through comprehensive experimentation. Subsequently, we will conduct an in-depth analysis and discussion based on the experimental results and consider potential ways to improve KAN’s performance in computer vision tasks.
@@ 4.1 Experiment 0: Preliminary Evaluation of KAN and CKAN on Core Vision Tasks
Result. As shown in Table 1, the experimental results indicate that models incorporating KAN and CKAN generally perform worse than the baseline models in classification and segmentation tasks, particularly the CKAN model. This result is unexpected, as the nonlinear characteristics of KAN and CKAN theoretically should enhance the model’s fitting capability. This observation motivates further investigation into why the nonlinear fitting abilities of KAN and CKAN have not been fully utilized in computer vision tasks.
@@
MobileNet First Layer CKAN 58.27
MobileNet Last Layer CKAN 58.64
MobileNet two CKANs 57.46
@@
UNet CKAN (Final CNN replaced) 59.13
@@ 4.2 Experiment 1: Effect of Dataset Size on KAN Performance
Results. The results, as shown in Table 3, indicate that CNN + KAN consistently outperforms the baseline CNN + MLP across all dataset sizes, demonstrating KAN’s stable fitting capability for this task. However, CKAN + CNN + MLP always performed worse than the baseline model, failing to exhibit the theoretical fitting advantage that CKAN should offer. Based on these experimental results, we can conclude that while KAN and CKAN theoretically possess stronger nonlinear fitting capabilities, in practice, CKAN has not effectively leveraged this advantage and performs worse than the traditional CNN + MLP. In contrast, KAN shows stable fitting performance, particularly as the dataset size increases, its performance improves. To investigate why CKAN failed to demonstrate its potential, we conducted Experiment 2.
@@
4.3 Experiment 2:
Based on the results of Experiment 1, we hypothesize that this may be due to the higher complexity of input in vision tasks, where image data is more likely to contain noise and outliers. The increased fitting capability may, therefore, lead to lower accuracy on the test set. To further investigate this issue, we designed Experiment 2, in which we intentionally introduce label noise to assess the performance of CKAN and KAN in noisy environments, with the aim of evaluating their robustness to noise in vision tasks.
@@ 4.3 Experiment 2:
Results. As shown in Table 4, as the noise level increases, the accuracy of all models decreases, with CKAN+CNN+MLP and CNN+KAN showing a more pronounced decline compared to the baseline CNN model. Although the initial performance of the KAN and CKAN models is slightly better than the baseline under low noise conditions, they exhibit significantly higher sensitivity to noise at higher noise levels. This indicates that KAN and CKAN have poor robustness, and their strong fitting capabilities do not seem suitable for handling high-complexity image inputs in computer vision tasks. Instead, they may lead to overfitting, thereby reducing the models’ performance in real-world applications.
@@ 4.4 Experiment 3: Impact of Regularization on the Generalization of KAN and CKAN
Results. As shown in Tables 5 and 6, we observed that with moderate L1 regularization, models incorporating KAN outperformed the baseline CNN model in accuracy, while models with CKAN consistently performed worse. This indicates that KAN’s fitting capability was better utilized, with reduced sensitivity to noise and outliers, highlighting the significant role of regularization in enhancing its generalization ability. In contrast, even with regularization, CKAN did not surpass the baseline model, and due to its high memory usage and slow computation speed, we concluded that the current CKAN structure has limited practical potential and is not a suitable alternative to CNNs. Based on these findings, we decided to abandon further exploration of CKAN and focus on improving KAN in the subsequent experiments, aiming to optimize its performance as the final layer of the model while enhancing its robustness.
@@ 4.5 Experiment 4: Enhanced Regularization and Segment Deactivation for Improved KAN Generalization
MobileNet + KAN + Smoothness Regularization Method + Segment Deactivation Technique 63.12
@@ 5 Conclusions
This study explores the application of Kolmogorov-Arnold Networks (KAN) and their variants in computer vision tasks. Despite KAN’s strong fitting capabilities in theory, our experimental results indicate that the high complexity of visual data and the prevalence of noise pose significant challenges to its generalization ability. This is particularly evident in the case of Convolutional KAN, which performs worse than traditional CNNs and is not suitable as a replacement for conventional CNNs. Similarly, directly replacing the NLP layer with KAN did not result in significant performance improvements across most tasks. To address these issues, we propose a smoothness regularization method and the ”Segment Deactivation” technique. These methods effectively reduce overfitting to noise and promote model smoothness, significantly enhancing the robustness of KAN. Ultimately, KAN achieves better performance when replacing the final NLP layer of the model, showing some potential for future applications.